Superior Markets, Inc., operates three stores in a large metropolitan area. A segmented absorption costing income statement for the company for the last quarter is given below:

For the Quarter Ended September 30 North Store South Store East Store Total $4,700,000 2,585,000 2,115,000 Sales Cost of goodNorth Store South Store East Store Total Administrative expenses: Store managers salaries General office salaries* InsuranceRequired 1 Required 3 Required 2 Required 4 Required 5 How much employee salaries will the company avoid if it closes the NorRequired 1 Required 3 Required 5 Required 2 Required 4 How much employment taxes will the company avoid if it closes the Nortequired 3 ed 1 Required 3 Required 4 Required 2 Required 5 What is the financial advantage (disadvantage) of closing the NortRequired 1 Required 2 Required 3 Required 5 Required 4 Assume that the North Stores floor space cant be subleased. However,
Answer #1

1.Salary of new employee = $13,200

Delivery crew = $5,700

Sales Salaries = $60,600

Store manager Salaries = $29,500- 14,200 = $15,300

General Office Salaries = $7,100

Employee Salaries Avoided per Quarter = $101,900

2.Employment Taxes Avoided = 101,900*15% = $15,285

3.Financial Advantage = costs avoided –margin lost

= Direct Advertising cost + Store Rent +2/3rd Insurance +Utilities + Salaries Avoided + Taxes Avoided- 360,000

= 68,000 + 86,000 + 12,600*2/3 + 26,250 + 101,900 + 15,285 – 360,000

= -$54,165 i.e. disadvantage

4.North store should not be closed, as there is a loss in closing the store

5.Financial Advantage of closing the north store = Gross Margin from Increased sales of East Store – Disadvantage calculated above

= (846,000/1,880,000)*940,000*1/4 – 54,165

=$51,585

Since there is net advantage, the store should be closed.